OpenStudy (anonymous):

7 cube root 5 -2 cuberoot 5

OpenStudy (anonymous):

(7-2)*\[\sqrt[3]{5}\]=5\[\sqrt[3]{5}\]

OpenStudy (anonymous):

\[5\sqrt{5}\]

OpenStudy (anonymous):

tie!!! lol

OpenStudy (anonymous):

i meant \[5\sqrt[3]{5}\]

OpenStudy (anonymous):

to late i got it wrong lmao

OpenStudy (anonymous):

wait. was it \[7\sqrt[3]{5}-2\sqrt[3]{5}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

5^5√3 was the answer

OpenStudy (anonymous):

so what is wrong with \[5\sqrt[3]{5}\]

OpenStudy (anonymous):

it must have asked something else.

OpenStudy (anonymous):

7 apples - 2 apples = 5 apples